Status of LHC installation & New Installation Schedule

Click here to load reader

download Status of LHC installation & New Installation Schedule

of 12

  • date post

    17-Jan-2016
  • Category

    Documents

  • view

    30
  • download

    0

Embed Size (px)

description

Status of LHC installation & New Installation Schedule. Planning issue. Sector 1-2, 2-3 & Point 3. LSS1R Q1 & Q3 reparation on surface, transport and lowering of low- β scheduled CW27-29 Arc 1-2 & DS Interconnection in progress, global advancement ~35% TI2 - PowerPoint PPT Presentation

Transcript of Status of LHC installation & New Installation Schedule

  • 19 June 2007Sylvain Weisz - LEMIC*Status of LHC installation&New Installation Schedule

    Sylvain Weisz - LEMIC

  • LSS1RQ1 & Q3 reparation on surface, transport and lowering of low- scheduled CW27-29Arc 1-2 & DSInterconnection in progress, global advancement ~35%TI2Magnets (MQI/MBI) and TED in place, beam-pipe installation in progress last LHC Arc cryo-magnets in UJ22 (MB2311/SSS652/MB3390) in placeLSS2LShort circuit tests in progressMKI not available, Alice ZDC installation CW24-25Low- repair & interconnection takes up to CW43 critical for completion of Sector 1-2LSS2RQ3.R2 on surface, DFBAD HCM in place, LCM available July 4th Short circuit tests completed last weekArc 2-3 & DSInterconnection in progress: ~70% global advancementQ11.R2 has a short and will be replaced (SSS ready CW32) impact on interconnect activityLSS3L/RDSLC: leaky compensators on thermal shield lines, should be replacedQ6.L3 and DFBMD available, installation pending completion of DSLCInstallation of last warm magnets postponed, pending DSLC installationModification of power cabling in UJ33 scheduled CW23-2719 June 2007Sylvain Weisz - LEMIC*Sector 1-2, 2-3 & Point 3

    Sylvain Weisz - LEMIC

  • Low- repair: conceptual drawing to implementation Cartridge model for fixed point reinforcementCartridge installation in Q1.R1 (done in Bld. 181)

    Sylvain Weisz - LEMIC

  • LSS1RQ1 & Q3 reparation on surface, transport and lowering of low- scheduled CW27-29Arc 1-2 & DSInterconnection in progress, global advancement ~35%TI2Magnets (MQI/MBI) and TED in place, beam-pipe installation in progress last LHC Arc cryo-magnets in UJ22 (MB2311/SSS652/MB3390) in placeLSS2LShort circuit tests in progressMKI not available, Alice ZDC installation CW24-25 ???Low- repair & interconnection takes up to CW43 critical for completion of Sector 1-2LSS2RQ3.R2 on surface, DFBAD HCM in place, LCM available July 4th Short circuit tests completed last weekArc 2-3 & DSInterconnection in progress: ~70% global advancementQ11.R2 has a short and will be replaced (SSS ready CW32) impact on interconnect activityLSS3L/RDSLC: leaky compensators on thermal shield lines, should be replacedQ6.L3 and DFBMD available, installation pending completion of DSLCInstallation of last warm magnets postponed, pending DSLC installationModification of power cabling in UJ33 scheduled CW23-2719 June 2007Sylvain Weisz - LEMIC*Sector 1-2, 2-3 & Point 3

    Sylvain Weisz - LEMIC

  • 19 June 2007Sylvain Weisz - LEMIC*Sector 3-4 & 4-5Arc 3-4 & DSInterconnection in progress, global advancement ~95%10 vacuum sub-sectors (out of 14) delivered to VAC, 7 leak tight and 3 under testReparation on-going on 500 series SSS (Q7-Q11.R3), need to open cold masses hold completion of interconnect in arc 3-4 LSS4L/RAll elements in place, NEG activation of straight section completed this week BGI detector damaged using wrong bake-out cycle!Klystron power tests in progress in UX45 Short to ground on MCB circuit in Q5.R4 reparation after 1st HW commissioning period of sector 4-5Arc 4-5 & DSInterconnection completed, pressure test passed successfully on May 11-12thHe flushing in progressCryo instrumentation checks, ELQA at warm and polarity checks in progressReplacement of 28 Velan valve puppets end of this week Cool-down of Sector 4-5 starts CW26 (next week)LSS5L Low-: DFBX/Q3/Q2/Q1.L5 back on surface for repair Cool-down 4-5 without low-Power test of warm magnets (D1.L5) completedFirst Roman Pots station in place

    Sylvain Weisz - LEMIC

  • 19 June 2007Sylvain Weisz - LEMICC*LSS5RReparation of low- fixed points scheduled CW27-33 critical path for completion of Sector 5-6Power test of D1.R5 completed, test of low- wire based alignment system in progressInterconnect Q4/Q5/Q6-DSLE and Q7-DFBAJ in progressArc 5-6 and DSInterconnection in progress, global advancement ~90%, 5 vacuum sub-sector completedLSS6LDFBAK: HCM, LCM and LEJL connection cryostat in placeSpecial configuration of DFBAK Proximity Equipment requiring extra ELQA (TP2) in placeMKD tests in progress, difficult co-activity with dump line installation dump line completion resumes after August 15thLSS6RInstallation of MKD and MKB tests in progressDFBAL HCM and LCM in place, ELQA (TP2) on-going with proximity equipment in placeArc 6-7 & DS:Interconnect in progress, global advancement ~55%LSS7L/RPreparation for warm magnet tests in progress (alignment smoothing done) - tests scheduled CW25-28Collimator supports in place, collimator installation starts CW29 after warm magnet tests Sector 5-6, 6-7 & Point 7

    Sylvain Weisz - LEMICC

  • Cryogenics conditions are now very stable below pointIce accumulation problems on current leads have been solved with the installation of dry air bagsD2, Q4 & Q5 (4.5K magnets) have been powered to their nominal current (6 kA)Main arc circuits (MB & QF/QD) and correctors chains have been qualified at coldQuench protection systems and associated heaters have been validatedTests of Powering Interlock Controller are almost complete Powering of the arc magnets is imminent But repairs in sector 7-8 (replacement of 1055, PIMs, bus-bar on SSS 500, etc) start early August Need 4 weeks for warm-up, power tests have to stop prematurely

    19 June 2007Sylvain Weisz - LEMIC*Temperature stabilisation and Power testsSector 7-8

    Sylvain Weisz - LEMIC

  • 19 June 2007Sylvain Weisz - LEMIC*Sector 8-1LSS8RLow-: HX reparation completed, interconnect (H pieces) completed, fixed point strenghtening (cartridge installation) CW25-31 critical path for completion of Sector 8-1Test of LHCb warm compensators (MBXW) completedY vacuum chamber available: D1-D2 vacuum installation in progressArc 8-1Interconnections completed Pressure test passed with success on June 9-10thLSS1LInterconnection of DFBLA to DSLA completedInstallation of D1-D2 vacuum chamber in progressLow-: HX reparation completed, interconnect (H pieces) completed, fixed point strenghtening (cartridge installation) CW25-31 critical path for completion of Sector 8-1

    Sylvain Weisz - LEMIC

  • Interconnections:Completed in Sectors 7-8, 8-1 & 4-5, pressure tests Ok92% of magnet to magnet electrical connections done, 51% of W closures completed Expect closure of sector 1-2 (arc & DS), incl. leak tests, by end of SeptemberUncertainties and delays from repair of Serie 500 bus-bar insulation

    Preparation for short circuit tests:Tests in progress in Point 2L (14/15 power station tested)Preparation for tests in Point 1R on scheduleTests in UJ33 will be re-done after modification of power cables routing

    Installation of warm beam pipes in long straight sections:Components procurement Ok, ~200m/wk installed and activatedAdvancement below expected rate, perturbed by:Area not available: LSS3L/R or UJ62-63 (for dump line)Co-activities: Point 6 (with kicker), Point 8R (with warm magnet tests)Uncertainty on equipment installation: collimators in LSS7L/R

    19 June 2007Sylvain Weisz - LEMIC*Installation Summary

    Sylvain Weisz - LEMIC

  • 25 May 2007Sylvain Weisz - ICC*Installation rate of LSS warm beam pipe Courtesy of Miguel Jimenez

    Sylvain Weisz - ICC

  • LHC Construction and InstallationGeneral co-ordination Schedule1223344556677881ConsolidationLHC project Document NLHC-PM-MS-0005.1 rev. 4.0

    CERN Dep./Group/Contractor Document N.TS/IC

    EDMS Document N.102509Interconnection of the continuous cryostatLeak tests of the last sub-sectorsInner Triplets repairs & interconnectionsGlobal pressure test &ConsolidationFlushingCool-downWarm upPowering Tests12 June 2007Prepared by Katy Foraz

    Sylvain Weisz - LEMIC

  • 19 June 2007Sylvain Weisz - LEMIC*New Installation Schedule - General FeaturesSector 4-5: Arc & DS ready cool-down without low-Sector 3-4: Delays with DSLC cool-down without DSLC Get experience with HW commissioning of 2 sectors in parallel Time window for fixes after 1st HW commissioning periodNeed for an extra warm-up/cool-down cycle

    Sector 7-8: Repairs scheduled Aug.-Sept. power tests stop early JulySector 8-1: Repair low- (8R & 1L) before cool-downNeed to warm-up sector 7-8 before completion of first power tests programCool-down and test low- as early as possible

    HW commissioning directly up to 7 TeVNeed to commission many sectors in parallelGet ready for physics by summer 2008

    Sylvain Weisz - LEMIC

    *********